Calculate number of neutrons present in 12 ×10^25 atoms of oxygen
.

Grade:11

2 Answers

Arun
4 years ago
Atomic no. is the number of electrons = no. of protons present in an atom.
Atomic Mass is the sum of the number of protons and neutrons present in an atom.
Let 
Atomic number = P = 8
Let Atomic mass = N + P = 17
N = 17 - 8
N = 9
Therefore there are 9 neutrons ina single atom of oxygen.
Therefore total no. of neutrons = 9 × 12 × 10^25 = 108 × 10^25 = 1.08 × 10^27
Answer
1.08 × 10^27 neutrons.
 
 
Khimraj
4 years ago
Atomic no. is the number of electrons = no. of protons present in an atom.
Atomic Mass is the sum of the number of protons and neutrons present in an atom.
Let 
Atomic number = P = 8
Let Atomic mass = N + P = 16
N = 16 - 8
N = 8
Therefore there are 8 neutrons ina single atom of oxygen.
Therefore total no. of neutrons = 8 × 12 × 10^25 = 86 × 10^25 = 9.6 × 10^26